OpenStudy (anonymous):

What is the product in lowest terms? -6/11 * 3/4 A. -9/22 B. -6/13 C. 9/22 D. 6/13

OpenStudy (nincompoop):

multiply the numerator and the same with denominator \[\frac{ a }{ c }\times \frac{ b }{ d }=\frac{ ab }{ cd }\]

OpenStudy (nincompoop):

then look for common factor to reduce it to lowest term

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

\[\frac{ 6 }{ 11 } \times \frac{ 3 }{ 4 } = \frac{ 33 }{ 24 }\]

OpenStudy (anonymous):

sORRY IM KIND OF CONFUSED RIGHT NOW....

OpenStudy (mathstudent55):

To multiply two fractions together, you multiply the numerators together, and that becomes the new numerator. Then you multiply the denominators together, and that becomes the new denominator. In this case, you are multiplying a positive number by a negative number, so the answer is negative.

OpenStudy (mathstudent55):

\(-\dfrac{6}{11} \times \dfrac{3}{4} \) \(= - \dfrac{6 \times 3}{11 \times 4} \) \(=- \dfrac{18}{44} \) Now we need to reduce the fraction. We need to find the largest number that divides evenly into both 18 and 44. This number is called the greatest common factor (GCF). In our case, the GCF is 2. We divide the numerator and denominator by 2 to reduce the fraction: \(= -\dfrac{9}{22} \) That is the final answer.

OpenStudy (anonymous):

sorry I was cooking thanks btw

OpenStudy (mathstudent55):

You're welcome.